OpenStudy (anonymous):

1/4 + 1/3 + 1/6 = the answer is 9/12.. but my question is how is it 9/12? i forgot how to do this can someone show me..explain please thanks :D

OpenStudy (anonymous):

Turn 1/4, 1/3, and 1/6 to have the same denominator: 1/4 = 3/12 (multiply top and bottom by 3) 1/3 = 4/12 (multiply top and bottom by 4) 1/6 = 2/12 (multiply top and bottom by 2) 3/12 + 4/12 + 2/12 = ?

OpenStudy (anonymous):

oh ok i got it now thanks! :D